OpenStudy (anonymous):

P=2(I+W) solve for w

OpenStudy (amistre64):

there is no w to solve for

OpenStudy (anonymous):

yes its in the brackets

OpenStudy (amistre64):

nope, all i see is P l and W. there is no w

myininaya (myininaya):

math is case sensitive

OpenStudy (anonymous):

that is the w

myininaya (myininaya):

its everything sensitive really

OpenStudy (anonymous):

come again..

OpenStudy (amistre64):

considering the obvious typo; how would you go about getting w all alone?

OpenStudy (anonymous):

we would be solving it, so it'd be w=blahblahblah

OpenStudy (anonymous):

idk thats why i asked for help

OpenStudy (amistre64):

good .... id start at least by dividng off that 2

OpenStudy (amistre64):

P = 2 (I+w) /2 /2 ---------- P/2 = l + w ; now what do you see we can do?

OpenStudy (anonymous):

w=p/2 -l?

OpenStudy (amistre64):

thats correct :) good job

OpenStudy (anonymous):

weoooo!

OpenStudy (amistre64):

\[w = \frac{P}{2}-l\]

OpenStudy (anonymous):

i have anothhhhhhhher question